Λ -DOUBLING IN HIGH ANGULAR MOMENTUM STATES: THE PURE ROTATIONAL SPECTRUM OF CoF (X^{3}Φ_{i})

\underline{M. A. FLORY}, P. M. SHERIDAN, M. A. BREWSTER, S. K. McLAMARRAH, L. M. ZIURYS, Dept. of Chemistry, Dept. of Astronomy, Steward Obsevatory, Univ. of Arizona, Tucson, AZ 85721; J. M. BROWN, Physical and Theoretical Chemistry Laboratory, Oxford University, Oxford OX1 3QZ, United Kingdom; and T. C. STEIMLE, Dept. of Chemistry and Biochemistry, Arizona State Univ., Tempe, AZ 85287.

The pure rotational spectrum of CoF (X ^{3}Φ_{i}) has been recorded in the frequency range 270-650 GHz - the first highresolution study to include all three spin components (Ω = 4, 3, and 2). CoF was created by reacting cobalt vapor with a mixture of 10% F_{2} in He. Fourteen rotational transitions were recorded. Λ-doubling was observed in both the Ω = 3 (5 MHz separation) and Ω = 2 (100 MHz) spin components, an unexpected result for a Φ state. In addition, the spectrum is further complicated by the presence of hyperfine interactions arising from both Co (I = 7/2) and F (I = 1/2) nuclei. The complete data set has been fit with a Hund case (a) Hamiltonian, and rotational, fine structure, Λ-doubling, and hyperfine parameters have been determined. The observation of Λ-doubling is in contrast to CoCl (X ^{3}Φ_{i}), where the effect was not observed.